Prove or disprove : for every real number x, |x − 8| + x ≥ 6.


1
Expert's answer
2021-11-10T15:47:35-0500

If "x>8" , then "|x-8|+x=x-8+x=2x-8>2\cdot 8-8=8" .

If "x\leq 8" , then "|x-8|+x=8-x+x=8" .

We have that "| x-8|+x\geq 8" .

Since "8>6" , it follows that "|x-8|+x\geq 6" for every real number "x" .